OpenStudy (anonymous):

Factor completely 3x2 + x + 7. (3x + 1)(x + 7) (3x + 7)(x + 1) Prime (3x + 4)(x + 3)

OpenStudy (anonymous):

\[here,\ 3x^2+x+7,\ we\ see\ that\ discriminant\\ D= b^2-4ac<0, where' \ b=1,\ a=3,\ c=7\\as\ D= 1^2-4.3.7=1-84=-83<0\\ \therefore\ its\ factors\ \ are\ [6x+(1+i \sqrt{83}][6x-(1-i \sqrt{83}]\\note:\ factors\ have\ been\ found\ out\ using\ zeros\ concept\ of \ polynomials\ and\\using\ Quadratic\ Formula.\]